Two bodies of equal charges each of 1C are separated in air by a distance of 1km. find the magnitude of the electric force between them.
1
Expert's answer
2020-06-08T10:15:12-0400

According to the Coulomb's law (https://en.wikipedia.org/wiki/Coulomb%27s_law) the force can be calculated as

"F = k\dfrac{q_1q_2}{\varepsilon r^2}," where "q_1" and "q_2" are the charges, "r" is the distance between charges, "k \approx 9\cdot10^9\,\mathrm{N\,m^2\,C^{-2}}" and "\varepsilon" is the permittivity, for air "\varepsilon = 1.000594 \approx 1."

Therefore,

"F = 9\cdot10^9\,\mathrm{N\,m^2\,C^{-2}}\cdot\dfrac{1\,\mathrm{C}\cdot1\,\mathrm{C}}{(1000\,\mathrm{m})^2} = 9\cdot10^3\,\mathrm{N}."